

********************************************************************
UARS WEEKLY STATUS REPORT - 29 DEC 2000
********************************************************************
***********
SOLSTICE
***********
Nothing received.
***********
SUSIM
***********
SUSIM performed daily solar scans of the sun during periods of
predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
AXIS 1 was activated on December 25. The PEM zenith spectrometers are
activate in daylight mode.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. Initial tests of telemetry transfers through
PEM2K:: have been successful; however, unexplained missing transfers have
occurred to PEM2K::. PEM Telemetry is runing in duplicate streams through
PEM:: and PEM2K::. We find that there are occasional files which transfer
from the CDHF to PEM::, but are missing from PEM2K::.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to operate normally. Work continues on our next software
delivery, which is to include updated wind and temperature recoveries, as well
as incorporating some new species (mesospheric ozone and stratospheric
aerosols). Happy holidays to all the UARS folks!!! Peace.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Daily processing of instrument Daily Activity Plans continues
with no problems through the holdiday season. Instrument
operations schedule is now complete to 9 February 2001.
***********
CDHF
***********
NRT, Q/L, and Level 0 data will be unavailable from 6:00 PM (EST) 12/30/00
through 8:00PM (EST) 12/31/00.
- Data created in the month of September 2000 has been revealed at the
DAAC.
- DAAC has concerns regarding the MLS V5 data that were transferred to
the DAAC Tables. These concerns continue to be addressed and resolved.
- WINDII Version 11 created in March, April, and May, 1999 has been loaded
into the DAAC Transfer Tables. DAAC has been notified.
UARS CDHF REHOST:
- Need positive response from all PI End Users prior to Rehost transition.
- PI validation of output products and concerns are being addressed.
- PEM V3.32 (JOBS 1 and 2) have been tested on the Rehost System.
- Job 1 has failed in the very last step (step 6, VMAG_SPLIT) and error
is currently being investigated. Job 2 was successful.
***********
FOT
***********
REPORTING PERIOD
This report covers 16 December 2000 (Orbit 50640, GMT Day 351) through 22
December 2000 (Orbit 50745, GMT Day 357).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 354/0112 50686 "Power ON,ZEPS/HEPS FULL TIME"
HRDI 354/0112 50686 Microloads
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 and 1.3 degrees C respectively. Additional battery performance data
for each battery on 15 December 2000 and 15 December 2000 are provided in
Appendix D.
The clock error ranged from Ð10.9 to 10.5 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 31
¥ PEM = 01
¥ HRDI = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 00
Three (3) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/315 - 2287.496493
· Transponder B: 00/314 - 2287.498584
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 00/342 show the observatory data loss to be 300 hours, 01
minutes, 24.888 seconds. This shows a zero loss since the last report period.
This is a 0.39 percent data loss, which equals a 99.61 percent data capture for
the mission.
· There was no loss for this week.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
This week the FOT uplinked new FDF calculated calibrations for both TAM1 and
TAM2. The calibration for TAM2 did not noticeable improve its performance,
therefore TAM1 was put into monitor mode. TAM1 has shown improved performance
over TAM2 due to its more orthogonal mounting, and will continue to be used
next week.
A new flight software patch to uplink a non-orthogonal matrix instead of a
quaternion is in its development stage.
MEETINGS
The FOT attended a TAM Attitude Meeting on 20 December 2000.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for December 31, 2000 (DOY 366). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 22 December 2000
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
50645 351/0652 00-728 NBTR NOOPS D-351
50651 351/1712 00-728 NBTR NOOPS D-351
50652 351/1810 00-729 NBTR NOOPS D-352
50660 352/0653 00-729 NBTR NOOPS D-352
50667 352/1811 00-730 NBTR NOOPS D-353
50674 353/0513 00-731 OBC Clk Adj 0.00 to 20.6 msec/day
50675 353/0653 00-730 NBTR NOOPS D-353
50681 353/1632 00-732 OBC/TAM MAGMOD UPDATE
50681 353/1632 00-733 OBC/TAM TAM ALIGNMENT UPDATE
50681 353/1714 00-730 NBTR NOOPS D-353
50681 353/1811 00-734 NBTR NOOPS D-354
50686 354/0112 00-735 PEM "Power ON,ZEPS/HEPS FULL TIME"
50686 354/0112 00-736 HRDI Microloads
50689 354/0654 00-734 NBTR NOOPS D-354
50692 354/1129 00-737 OBC Clk Adj 20.6 to 0.00 msec/day
50696 354/1715 00-738 OBC/TAM TAM1 ALIGNMENT UPDATES
50696 354/1715 00-734 NBTR NOOPS D-354
50697 354/1813 00-739 NBTR NOOPS D-355
50698 354/2035 00-740 ACS TAM1 QUARTERNION
50698 354/2035 00-741 ACS TAM1 MONITOR MODE
50705 355/0656 00-739 NBTR NOOPS D-355
50711 355/1634 00-742 ACS TAM TO UPDATE MODE
50712 355/1814 00-743 NBTR NOOPS D-356
50719 356/0615 00-743 NBTR NOOPS D-356
50725 356/1537 00-744 ACS TAM ERROR VARIANCE
50726 356/1635 00-743 NBTR NOOPS D-356
50726 356/1718 00-745 ACS FSS TP UPDATE MODE/TAM MONITOR
50727 356/1815 00-747 NBTR NOOPS D-357
50730 356/2334 00-746 OBC STATBUF RESET
50734 357/0616 00-747 NBTR NOOPS D-357
50739 357/1356 00-749 OBC Table 37 Up-date
50741 357/1636 00-747 NBTR NOOPS D-357
50742 357/1817 00-748 NBTR NOOPS D-358
__________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 22 December 2000
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 22 December 2000
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
50674 353/0514 Clk Adj 0.00 to 20.6 msec/day
50692 354/1132 Clk Adj 20.6 to 0.00 msec/day
__________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 22 December 2000
15 Dec 2000 - GMT Day 350 Beta = 33.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ---------- ------ ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.2 0.000 0.0
2 +16.8/-44.8 +5.79/+4.51 +28.8/-15.2 25.0 1.050 21.0
3 +22.4/-112.0 +4.51/+3.23 +28.8/-15.2 25.0 1.040 21.0
22 Dec 2000 - GMT Day 357 Beta = 27.6 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 ,+728.0/+728.0 -3.05/-4.31 +0.0/+0.0 14.2 0.000 0.0
2 +39.2/-28.0 +5.79/+4.83 +30.4/-14.4 24.5 1.040 21.3
3 +22.4/-112.0 +5.15/+3.87 +29.6/-15.2 24.5 1.040 21.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---------- ----------
no changes V/T 5
_________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 15 December 2000
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-047 FOT see list Nov 00 14170 Generic Late Acq.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
00-053 FOT 50168 319/1113 23157 Late Acquisition
00-054 FOT 50280 326/2147 23187 RESET PTP /2:39 32K
00-056 FOT 50584 347/0502 23302 Premature LOS.
SEVNTFW00/51